72 6) Find the derivative of f [2 - (2x + 1) using the product rule. a) 0 2 + + 1 b) 4 + 1 c) O 2 - 1 х d) - 2
Answer #1

\text{ Let }

f(x) =\left [ 2-\frac{1}{x} \right ]\left ( 2x+1 \right )

Calculating the derivative of f(0)

f'(x) =\frac{\mathrm{d} }{\mathrm{d} x}\left (\left [ 2-\frac{1}{x} \right ]\left ( 2x+1 \right ) \right )

=> f'(x) =\left (\frac{\mathrm{d} }{\mathrm{d} x}\left [ 2-\frac{1}{x} \right ] \right ) \left ( 2x+1 \right ) +\left [ 2-\frac{1}{x} \right ]\left ( \frac{\mathrm{d} }{\mathrm{d} x} \left ( 2x+1 \right ) \right )

\left ( \because \text{ Product rule, }\:\:\:\frac{\mathrm{d} }{\mathrm{d} x} (uv)=\left ( \frac{\mathrm{d} u}{\mathrm{d} x} \right )v+u\left ( \frac{\mathrm{d} v}{\mathrm{d} x} \right )\right )

=> f'(x) =\left (\frac{\mathrm{d} }{\mathrm{d} x}(2)-\frac{\mathrm{d} }{\mathrm{d} x} \left (\frac{1}{x} \right )\right ) \left ( 2x+1 \right ) +\left [ 2-\frac{1}{x} \right ] \left ( \frac{\mathrm{d} }{\mathrm{d} x} ( 2x)+\frac{\mathrm{d} }{\mathrm{d} x}(1) \right )

=> f'(x) =\left (0- \left (-\frac{1}{x^2} \right )\right ) \left ( 2x+1 \right ) +\left [ 2-\frac{1}{x} \right ] \left ( 2(1)+0 \right )

=> f'(x) =\left (\frac{1}{x^2} \right ) \left ( 2x+1 \right ) +\left [ 2-\frac{1}{x} \right ] \left ( 2 \right )

=> f'(x) =\frac{2x}{x^2} +\frac{1}{x^2} + 4-\frac{2}{x}

=> f'(x) =\frac{2}{x} +\frac{1}{x^2} + 4-\frac{2}{x}

=> f'(x) =\frac{2}{x} +\frac{1}{x^2} + 4-\frac{2}{x}

=> f'(x) =\frac{1}{x^2} + 4

Hence,

f'(x) =4+ \frac{1}{x^2}

\text{ Option b is correct.}